OpenStudy (tw101):

Help with area of a triangle with vertices: Find area of a triangle with vertices (-8,10),(6,17), and (2,-4) I have no idea where to even start. Thanks!

OpenStudy (freckles):

use heron's formula

OpenStudy (tw101):

I haven't learned that yet...

OpenStudy (tw101):

what is it?

OpenStudy (freckles):

it is a formula to find the area of a triangle

OpenStudy (freckles):

using the lengths of the sides

OpenStudy (mathmale):

That's a great approach, but one I've used before. An alternative approach would be to graph the given vertices of the triangle and then use either the distance formula or the Pyth. Theorem to determine the length of the base and the height of this triangle. Then, Area A = (1/2)*base*height

OpenStudy (tw101):

Okay. Do you use the distance formula to figure out the distances between the points?

OpenStudy (tw101):

@mathmale

OpenStudy (freckles):

yes

OpenStudy (tw101):

Okay

OpenStudy (tw101):

Is that all you need to do then?

OpenStudy (tw101):

Sorry i just realized he said distance or pyth. theorem already haha

OpenStudy (sweetnikkie072):

find side using distnce formula d= √(x2-x1)^2+(y2-y1)^2 use triangle's area= 1/2*base*height using the above found sides

OpenStudy (tw101):

Okay thanks you guys! that helps

OpenStudy (sweetnikkie072):

you can use pythagoras theorem only when trianle is 90º aka right angled triangle

OpenStudy (tw101):

Gotcha. I wish i could medal all of you guys!

OpenStudy (freckles):